1. A $4,000 , 12% bond will be redeemed on
May 1,2008. Interest is payable semiannually
on May 1 and November 1.
Find the purchase price of the bond if the date of purchase is May 1, 1998 and
the yield rate is 10% compounded semi-annually.
1
Expert's answer
2020-06-04T10:42:12-0400

"CF=\frac{4000\times0.12}{2}=240"

N=4000


rates from the table of discounts and multiples


"P=CF\times(FM4(\frac{r}{2};2\times n)+N\times(FM2(\frac{r}{2};2\times n)=240\times(FM4(5;20)+N\times(FM2(5;20)=240\times12.462+4000\times0.377=2 990,88+1508=4498.88"